Ultraviolet light of wavelengths 800 Å when allowed to fall on hydrogen atoms in their ground state is found to liberate electrons with kinetic energy 1.8 eV  and 4.0 eV respectively. Find the value of Planck’s constant.

Grade:11

1 Answers

Kevin Nash
9 years ago
hc / λ1 – hc / λ0 = K.E. 1 ….(i)
And hc / λ2 – hc / λ0 = K.E.2 ….(ii)
⇒ hc / λ1 - hc / λ2 = K.E.1 – K.E2
⇒ hc [λ2 - λ1 / λ1 λ2] =K.E.1 – K.E2
∴ h = (K.E. – K.E.­21 λ2 / c (λ2 – λ1)
= (1.8 – 4) x 1. 6 x 10-19 x 800 x 10-10 x 700 x 10-10 / 3 x 108 x (700 - 800) x 10-10
= 6.6 x 1034 J.s.